I need to create a real life situation that fits the equation (x+3)(x-5)=0 and express the situation as the same equation.
Answer by Mathtut(3670)

Sam is 2 years younger than Sally. The product of their ages is 15, how old is Sam?
Sally is x, Sam is x-2
x(x-2)=15
x^2-x=15
x^2-x-15=0
(x + 3)(x - 5) = 0